# State Entropy Regularization for Robust Reinforcement Learning

### v2

@@ A.2 Proof of results in Table 1
Next, we prove the limiting case when $\alpha\to\infty$. Using l’hospital’s rule we get:
@@
We finally prove the limiting case for $\alpha\to 0$. Again using l’hospital’s rule:
@@
[…]
@@ A.4.1 Policy entropy lower bound
We consider the same lower bound as in equation 28 and the same definition of $\Delta r^{\prime}$ as in 29. The uncertainty set ~$\tilde{{\mathcal{R}}}^{\pi}(\epsilon,\alpha)$ corresponding to state-action entropy regularization is (see Table 1):
@@
[…]
@@
Fix some $M>0$, $\alpha>0$, and confidence level $10<\beta<$ 1. In this proof we’ll denote the total discounted return by $G$. We construct an MDP (illustrated in Figure 6) that induces arbitrarily poor risk-sensitive performance under occupancy entropy regularization.
@@
We now show explicitly that exists $R_{3},R_{4},p$ s.t. the two conditions above hold. We focus on the second condition, and use the dual representation of CVaR [2]: $\mathrm{CVaR}_{\beta}^{\pi_{R}^{*}}[G]=\min_{Q\in\mathcal{U}}\mathbb{E}_{G\sim Q}[G]$ where $\mathcal{U}=\{Q\ll P\mid\forall G:\frac{Q(G)}{P(G)}\in(0,1/\beta)\}$ and $P$ is the return distribution for $\pi^{*}_{R}$.
Since $01>\beta>$ 0, we can choose $01>p>$ 0 s.t. $(1-q)p+q\geq\beta$. We construct a probability measure $Q\in\mathcal{U}$ s.t. $Q(R_{3})=\frac{p(1-q)}{\beta}$ and $Q(R_{1})=1-Q(R_{3})$. Note that from the construction of $p$ we have that $Q(R_{3})<1$ and $1-Q(R_{3})=Q(R_{1})\leq\frac{P(R_{1})}{\beta}=\frac{q}{\beta}$ and indeed $Q\in\mathcal{U}$. We can now analyze the risk averse performance of $\pi^{*}_{R}$:
@@
Figure 6: Counterexample used in the proof of Theorem 4.2. The initial state is $s_{0}$. All actions are deterministic except for action $a$ taken from state $s_{2}$, which leads stochastically to $s_{3}$ with probability $1-p$ and to $s_{4}$ with probability $p$. Self-loops indicate absorbing states with associated returns.
@@ Appendix B Experiments
The code for our experiments is available at https://anonymous.4open.science/r/State_entropy_robust_rl-3810.
@@ B.1 State entropy algorithm
k-nearest neighbor entropy estimator. Let $X$ be a random variable with a probability density function $p$ whose support is a set $\mathcal{X}\subset\mathbb{R}^{q}$ Then its differential entropy is given as $\mathcal{H}(X)=-\mathbb{E}_{x\sim p(x)}[\log(p(x))]$ When the distribution $p$ is not available, this quantity can be estimated given N i.i.d realizations of $\{x_{i}\}_{i=1}^{N}$ However, since it is difficult to estimate $p$ with high-dimensional data, particle-based k-nearest neighbors (k-NN) entropy estimator [51] can be employed:
@@
Where $x_{i}^{k\text{-NN}}$ is the k-NN of $x_{i}$ within a set $\{x_{i}\}_{i=1}^{N}$,$C_{k}=log(k)-\Psi(k)$ is a bias correction term,$\Psi$ the digamma function,$\Gamma$ the gamma function, q the dimension of x, $\hat{\pi}\approx 3.14159$ , and the transition from (1) to (2) always holds for q > 0.
@@
$r^{i}(s_{i})=\log(||y_{i}-y^{k-nn}_{i}||+1)$ (3)
where $y_{i}$ can be either $s_{i}$, a latent encoding of $s_{i}$ or selected features of $s_{i}$ The total reward is:
@@
Temperature warm-up. We find that it helps performance to start training with a high temperature $\beta_{start}$ and anneal it through training. This is common in prior work [49], notably instead of exponential decay we use cosine decay from $\beta_{start}$ to $\beta>0$
@@ B.2 Minigrid experiments
The experiment is built upon [25] open code base, which is released under the MIT License . [^4] Available at https://github.com/kingdy2002/VCSE. Base algorithm is A2C. As the point of the expirement isnt handling partial observability, we remove potential bias by instead of using a random encoder, the features used for entropy estimation were the x,y position of the agent + one hot encoding of its direction. The nominal environment is an 8 by 8 empty grid with a sparse reward, 1 if goal is reached, 0 otherwise. action space is (1) take a step in current direction, (2) turn right (3) turn left.
@@ B.3 Pusher experiment
The code for our experiments is based on the CleanRL codebase [19], which is released under the MIT License  [^5] https://github.com/vwxyzjn/cleanrl We use PPO with a recurrent network for all baselines. Note that it is known that to maximize state entropy over finite number of rollouts, a non-Markovian policy is necessary. State entropy is calculated over the puck’s x,y positions. The obstacle is composed of three axis-aligned blocks with width 3cm, centered at (0.52, -0.24), (0.55, -0.21) and (0.58, -0.18). These positions are roughly along the perpendicular bisector of the line between the puck’s initial position and the goal. We used 200 episodes of length 100 for evaluating each method. As in [12], to decrease the variance in the results, we fixed the initial state of each episode:
$qpos$ =[0.,0.,0.,0.,0.,0.,0.,0.3,0.2,0.,0.=[0.,0.,0.,0.,0.,0.,0.,-0.3,0.2,0.,0.]= [ 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , - 0.3 , 0.2 , 0 . , 0 . ]
$qvel$ =[0.,0.,0.,0.,0.,0.,0.,0.,0.,0.,0.=[0.,0.,0.,0.,0.,0.,0.,0.,0.,0.,0.]= [ 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , 0 . , 0 . ]
@@
B.4 Interplay between State and Policy Entropy Regularization

### v3

@@ A.2 Proof of results in Table 1
Next, we prove the limiting case when $\alpha\to\infty$. Using L’Hôpital’s rule we get:
@@
We finally prove the limiting case for $\alpha\to 0$. Again using L’Hôpital’s rule:
@@
$\log\mathbb{E}_{\rho^{\pi}_{\tilde{P}}}[r(s,a)]$ $=\log\mathbb{E}_{\rho^{\pi}_{{P}}}\left[\frac{\rho^{\pi}_{\tilde{P}}}{\rho^{\pi}_{P}}r(s,a)\right]\qquad\text{(Importance sampling)}$ (26)
$\geq\mathbb{E}_{\rho^{\pi}_{{P}}}\left[\log(r(s,a))\right]+\mathbb{E}_{\rho^{\pi}_{{P}}}\left[\log\frac{\rho^{\pi}_{\tilde{P}}}{\rho^{\pi}_{P}}\right]\qquad\text{(Jensen)}$ (27)
@@ A.4.1 Policy entropy lower bound
We consider the same lower bound as in equation 28 and the same definition of $\Delta r^{\prime}$ as in 29. The uncertainty set $\tilde{{\mathcal{R}}}^{\pi}(\epsilon,\alpha)$ corresponding to state-action entropy regularization is (see Table 1):
@@
Let $\Omega\not\equiv 0$ a function. Assume towards contradiction that exists a function $\tilde{\mathcal{P}}(P)$ that satisfies $\min_{\tilde{P}\in\tilde{\mathcal{P}}}\mathbb{E}_{\rho^{\pi}_{\tilde{P}}}[r]=\mathbb{E}_{\rho^{\pi}_{P}}[r]+\Omega(\pi,P)$. We consider an MDP with reward vector $R\equiv c$ where $c\in\mathbb{R}$ is some constant. For all $\pi,P$ we have:
@@
Before the proof, we state the formal definition of the Conditional Value-at-Risk (CVaR). Let the return be defined as $R(\tau):=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$ where the trajectory $\tau=(s_{0},a_{0},r_{0},\ldots)$ is sampled from $p^{\pi}_{P}(\tau):=\mathbb{P}[s_{0},a_{0},r_{0},\ldots|\pi,P]$. Then, for a given confidence level $\beta\in(0,1)$ the CVaR of the return under policy $\pi$ is defined as:
where $(x)^{+}:=\max(x,0)$.
[…]
Figure 7: Counterexample used in the proof of Theorem 4.2. The initial state is $s_{0}$. All actions are deterministic except for action $a$ taken from state $s_{2}$, which leads stochastically to $s_{3}$ with probability 1p1-p and to $s_{4}$ with probability pp. Self-loops indicate absorbing states with associated returns.
@@
Fix some $M>0$, $\alpha>0$, and confidence level $0<\beta<1$. In this proof we’ll denote the total discounted return by $G$. We construct an MDP (illustrated in Figure 7) that induces arbitrarily poor risk-sensitive performance under occupancy entropy regularization.
@@
We now show explicitly that exists $R_{3},R_{4},p$ s.t. the two conditions above hold. We focus on the second condition, and use the dual representation of CVaR [Ang2017]: $\mathrm{CVaR}_{\beta}^{\pi_{R}^{*}}[G]=\min_{Q\in\mathcal{U}}\mathbb{E}_{G\sim Q}[G]$ where $\mathcal{U}=\{Q\ll P\mid\forall G:\frac{Q(G)}{P(G)}\in(0,1/\beta)\}$ and $P$ is the return distribution for $\pi^{*}_{R}$.
Since $1>\beta>0$, we can choose $1>p>0$ s.t. $(1-q)p+q\geq\beta$. We construct a probability measure $Q\in\mathcal{U}$ s.t. $Q(R_{3})=\frac{p(1-q)}{\beta}$ and $Q(R_{1})=1-Q(R_{3})$. Note that from the construction of $p$ we have that $Q(R_{3})<1$ and $1-Q(R_{3})=Q(R_{1})\leq\frac{P(R_{1})}{\beta}=\frac{q}{\beta}$ and indeed $Q\in\mathcal{U}$. We can now analyze the risk averse performance of $\pi^{*}_{R}$:
@@
A.7 Robustness under estimation errors
In Section 3, we provided robustness guarantees for the reward uncertainty (Theorem 3.1) and kernel uncertainty (Theorem 3.3) of a policy maximizing the state-entropy regularized objective. In practice, however, the regularization term is not computed with the true state entropy but a sample-based estimate. It is then not obvious than similar guarantees hold under estimation error of the state entropy. In the following, we show how the guarantees generalize to the sample-based regime following a straightforward concentration argument on the state distribution. In the spirit of Section 3, we consider the entropy of a finite state space, while we refer to Section 4.3 for an informal discussion on the sample sensitivity in continuous settings, in which more sophisticated estimators are employed.
Corollary A.1.
Let $\hat{d}$ an empirical estimate of $d^{\pi}_{P}$ obtained with $n$ sampled trajectories, for which it holds $\|\hat{d}-d^{\pi}_{P}\|_{1}\leq 1/2$. For $\delta\in(0,1)$ and any policy $\pi$, the following weak duality holds with probability $1-\delta$
$\min_{\tilde{r}\in\tilde{\mathcal{R}}}\mathbb{E}_{\rho^{\pi}}[\tilde{r}]\geq\mathbb{E}_{\rho^{\pi}}[r]+\alpha\Bigg({\mathcal{H}}_{{\mathcal{S}}}(\hat{d})-\sqrt{\frac{2S\log(2/\delta)}{n}}\log\sqrt{\frac{Sn}{\log(2/\delta)}}-\log(S)\Bigg)-\epsilon,$
where the reward uncertainty set is $\tilde{\mathcal{R}}^{\pi}(\epsilon,\alpha):=\left\{\tilde{r}\mid\textsc{lse}_{{\mathcal{S}}}(\frac{r^{\pi}-\tilde{r}^{\pi}}{\alpha})\leq\frac{\epsilon}{\alpha}+\log(S)\right\}$.
Proof.
@@ A.7 Robustness under estimation errors
$|{\mathcal{H}}_{{\mathcal{S}}}(\hat{d})-{\mathcal{H}}_{{\mathcal{S}}}|$ $\leq\kappa\log S+\kappa\log\frac{1}{\kappa}+(1-\kappa)\log\frac{1}{1-\kappa}$ (45)
$\leq 2\kappa\log\frac{S}{\kappa}$ (46)
$\leq\sqrt{\frac{2S\log(2/\delta)}{n}}\log\sqrt{\frac{Sn}{\log(2/\delta)}}\qquad\qquad\text{[with probability }1-\delta\text{]}$ (47)
where the first inequality follows from continuity bounds of the entropy [see, winter2016tight], the second inequality follows from the assumption $\kappa\leq 1/2$, the last inequality invokes concentration of the empirical distribution [weissman2003inequalities]. Finally, the result is straightforward by plugging equation 47 into the result of Theorem 3.1. ∎
An analogous result for kernel uncertainty is as follows.
Corollary A.2.
Let $\hat{d}$ an empirical estimate of $d^{\pi}_{P}$ obtained with $n$ sampled trajectories, for which it holds $\|\hat{d}-d^{\pi}_{P}\|_{1}\leq 1/2$. For any $\delta\in(0,1)$, policy $\pi$, and $\alpha>0$, the following weak duality holds with probability $1-\delta$
$\min_{\tilde{{\mathcal{P}}}^{\pi}(\epsilon)}\mathbb{E}_{\rho^{\pi}_{\tilde{P}}}[r]\geq\mathrm{exp}\Bigg(\mathbb{E}_{\rho^{\pi}_{P}}[\log(r)]+\alpha\bigg({\mathcal{H}}_{{\mathcal{S}}}(\hat{d})-\sqrt{\frac{2S\log(2/\delta)}{n}}\log\sqrt{\frac{Sn}{\log(2/\delta)}}-\log(S)\bigg)-\epsilon\Bigg),$
where
$\tilde{{\mathcal{P}}}^{\pi}(\epsilon,\alpha):=\bigg\{\tilde{P}\mid\alpha\log\bigg(\frac{1}{{\mathcal{S}}}\sum_{s\in{\mathcal{S}}}\bigg(\frac{d^{\pi}_{P}(s)}{d^{\pi}_{\tilde{P}}(s)}\bigg)^{\frac{1}{\alpha}}\bigg)\leq\epsilon\bigg\}$
Proof.
The proof is straightforward by plugging equation 47 into the result of Theorem 3.3. ∎
@@ Appendix B Experiments
The code for our experiments is available at https://github.com/JonathanAshlag/State_entropy_robust_rl.
@@ B.1 State entropy algorithm
k-nearest neighbor entropy estimator. Let $X$ be a random variable with a probability density function $p$ whose support is a set $\mathcal{X}\subset\mathbb{R}^{q}$ Then its differential entropy is given as $\mathcal{H}(X)=-\mathbb{E}_{x\sim p(x)}[\log(p(x))]$ When the distribution $p$ is not available, this quantity can be estimated given N i.i.d realizations of $\{x_{i}\}_{i=1}^{N}$ . However, since it is difficult to estimate $p$ with high-dimensional data, particle-based k-nearest neighbors (k-NN) entropy estimator [singh2003nearest] can be employed:
@@
[…]
@@
[…]
where $y_{i}$ can be either $s_{i}$, a latent encoding of $s_{i}$ or selected features of $s_{i}$ . The total reward is:
@@
Temperature warm-up. We find that it helps performance to start training with a high temperature $\beta_{start}$ and anneal it through training. This is common in prior work [seo2021state], notably instead of exponential decay we use cosine decay from $\beta_{start}$ to $\beta>0$
@@ B.2 Minigrid experiments
The experiment is built upon [kim2024accelerating] open code base, which is released under the MIT License . [^4] Available at https://github.com/kingdy2002/VCSE. Base algorithm is A2C. As the point of the expirement isnt handling partial observability, we remove potential bias by instead of using a random encoder, the features used for entropy estimation were the x,y position of the agent + one hot encoding of its direction. The nominal environment is an 8 by 8 empty grid with a sparse reward, 1 if goal is reached, 0 otherwise. action space is (1) take a step in current direction, (2) turn right (3) turn left.
@@ B.3 Pusher experiment
The code for our experiments is based on the CleanRL codebase [huang2022cleanrl], which is released under the MIT License  [^5] https://github.com/vwxyzjn/cleanrl We use PPO with a recurrent network for all baselines. Note that it is known that to maximize state entropy over finite number of rollouts, a non-Markovian policy is necessary. State entropy is calculated over the puck’s x,y positions.
Transition kernel perturbation (wall): The obstacle is composed of three axis-aligned blocks with width 3cm, centered at (0.52, -0.24), (0.55, -0.21) and (0.58, -0.18). These positions are roughly along the perpendicular bisector of the line between the puck’s initial position and the goal. We used 200 episodes of length 100 for evaluating each method. As in [eysenbach2022maximum], to decrease the variance in the results, we fixed the initial state of each episode:
$qpos$ $=[0.,0.,0.,0.,0.,0.,0.,-0.3,0.2,0.,0.]$
$qvel$ $=[0.,0.,0.,0.,0.,0.,0.,0.,0.,0.,0.]$
Action robust baseline [Chen2019] We ran as a baseline both PRMDP and NRMDP. To ensure a fair comparison we implemented the following modifications:
  • Larger networks: We increased the width of all neural networks to 256 units.
  • hyperparameter sweep: we swept for exploration noise $\in[0.2,0.6,1]$, learning rate $\in[1e-3,5e-4,1e-4]$ and batch size $\in[256,512,1024]$
  • Check-pointing: we saved both the best performing checkpoint during training (evaluated in the nominal environment) and the last iteration checkpoint. We evaluated both of them in the perturbed environment and showed the better of the two.
Reward shift: The original goal was located at (0.66, -0.35). In the reward shift experiment, a new goal is randomly sampled uniformly from the circumference of a circle with radius r centered at the original goal location. For each radius in [0.1,0.15] we randomly sample 200 goals, evaluate once per goal, and report the average performance across all goals.
@@
B.4 Ant experiment
We train All policies on Mujoco’s "ant-v5". To emphasize the affects of each regularization we augment the environment to be less reward shaped in the following way: The episode does not truncate when the ant’s in an unhealthy position. Furthermore, there is no reward for being in a healthy position and the penalties on enduring contact forces and control cost are turned off. The reward in our setup is only the forward progress reward. State entropy is calculated over the height and orientations of the ant, encouraging the agent to learn to walk in diverse ways.
Transition kernel perturbation (vertical wall) During evaluation, we add an unobservable wall of a size of $1,\infty,0.3]$; the only way to pass the wall is to hurdle above it. Evaluation protocol is the same as in the pusher experiment.
B.5 Interplay between State and Policy Entropy Regularization